# FLARE: Fast Low-rank Attention Routing Engine

### v2

@@ 3.3 Spectral analysis
The matrix $W_{h}$ in Eq. 9 represents the attention weights between $N$ tokens, where $W_{h}]_{ij}$ quantifies how much token $j$ communicates to token $i$ within head $h$. Since $W_{h}$ has rank at most $M$, the model can capture at most $M$ independent global communication patterns per head. The eigenvalues of $W$ indicate the relative importance or energy of each latent dimension in forming the attention matrix $W$. We apply an algorithm to obtain the eigen decomposition of $W$ in $\mathcal{O}(M^{3}+M^{2}N)$ time compared to $\mathcal{O}(N^{3})$ for a dense communication matrix. The algorithm is predicated on computing the eigenspectra of an $M\times M$ matrix $JJ^{T}$ where $J\in\mathbb{R}^{M\times N}$ is chosen such that $J^{T}J$ is similar to $W$. The algorithm is detailed in Section B.1, and summarized in Algorithm 1. Figure 7 presents the $M$ nonzero eigenvalues of $W_{h}$ for an FLARE model trained on the elasticity benchmark problem with 972 points per input. The distinct spectra of the heads indicates that each head learns distinct attention patterns.
The eigenvalue analysis detailed in Section B.2 shows that while FLARE provides capacity for rank-$M$ attention, the model learns to use only a small fraction of this in early blocks indicating effective compression. In deeper blocks, more of the latent capacity is utilized, with diverse spectral profiles across heads, validating our design choice of independent head-wise projections.
@@ 4 Experiments
Model
Elasticity
Darcy
Airfoil
Pipe
DrivAerML-40k
LPBF
Vanilla Transformer
(Vaswani et al., 2017)
5.37
660k
4.38
660k
6.28
660k
$\sim$ $\sim$ $\sim$
PerceiverIO
(Jaegle et al., 2021a)
23.4
1.93m
21.5
1.93m
162
1.93m
7.14
1.93m
760
1.93m
56.3
1.93m
GNOT
(Hao et al., 2023)
13.3
4.87m
16.9
4.90m
103
4.90m
5.89
4.90m
115
4.87m
24.3
4.87m
LNO
(Wang & Wang, 2024a)
9.25
1.83m
7.64
762k
17.8
762k
8.10
762k
146
762k
24.7
762k
Transolver w/o conv
(Wu et al., 2024)
6.40
713k
18.6
713k
8.24
713k
4.87
713k
70.5
713k
20.4
713k
Transolver with conv
(Wu et al., 2024)
$\backslash$
5.94
2.8m
5.50
2.8m
3.90
2.8m
[…]
FLARE (ours)
3.38
592k
5.10
691k
4.28
691k
2.85
625k
60.8
691k
18.5
625k
@@ 4.1 PDE surrogate benchmarks
We consider a diverse set of benchmark datasets (Table 3) for regressing PDE solutions on point clouds spanning structured and unstructured grids with up to 50,000 points. Note that FLARE is mesh-agnostic, and operates solely on the input point cloud. The 2D elasticity, darcy, airfoil, and pipe benchmarks (Li et al., 2020; 2023b) cover a wide range of physical phenomena, and the 3D DrivAerML benchmark (Ashton et al., 2024) provides automotive aerodynamic simulations. Additional details of the dataset are presented in Section C.2. We also introduce a 3D field-prediction benchmark derived from laser powder bed fusion (LPBF) simulations, with diverse 3D-printed parts containing up to 50,000 grid points (see Appendix G).
@@
We compare FLARE with state-of-the-art PDE surrogates: generic attention models (vanilla Transformer (Vaswani et al., 2017), PerceiverIO (Jaegle et al., 2021a)); attention-based PDE surrogates (Transolver (Wu et al., 2024), LNO (Wang & Wang, 2024a)); and the neural operator GNOT (Hao et al., 2023). We exclude graph-based models because graph connectivity is unavailable for most problems. Transolver++ (Luo et al., 2025) is a concurrent work; although we tested the recently released official implementation with reported hyperparameters, preliminary results did not match those reported in the paper, so we do not include it in our comparisons. We follow the experimental setup of Transolver as it is the preeminent surrogate model and attempt to match its parameter count. Note that Transolver can be instantiated in two configurations: without convolution, where point-to-point communication relies solely on physics attention, and with convolution, where convolution layers are added to inject information from neighboring points when the input grid is structured. We evaluate these two configurations separately to isolate the impact of convolution versus physics attention. In our model, we choose not to employ any convolution layers and rely entirely on FLARE for token mixing. Because these PDE problems are relatively small (up to 50,000 points), we train all models in FP32. As shown in Figure 8, the vanilla Transformer is drastically slower than FLARE and Transolver on large point clouds; accordingly, we evaluate it only on problems up to $\sim 10{,}000$ points.
@@ 4.2 Field-prediction on million-point geometries
Although the benchmark problems in Section 4.1 represent a wide variety of PDE problems, they are relatively small compared to industrial use cases that demand PDE solutions on complex geometries with millions of grid points (Ashton et al., 2024). So far, attention-based surrogate models have not been able to scale to million-scale regression problems due to quadratic time and memory complexity, as illustrated in Figure 2. The flash attention (Dao et al., 2022) algorithm has alleviated the memory bottleneck thanks to online softmax computation; however, these methods remain impractical due to their long training times. Furthermore, SOTA models such as Transolver and LNO cannot be implemented with off-the-shelf fused attention algorithms like flash attention because of the need to explicitly materialize the projection weights.
@@
We demonstrate in Figure 4 that FLARE can scale to million-scale geometries by training on the DrivAerML dataset (Ashton et al., 2024) where each mesh is subsampled to contain one million points. These calculations are performed in mixed precision on a single Nvidia H100 80GB GPU provisioned through Google Cloud Platform. We note the clear trend in Figure 4 (left) that the error consistently decreases as we increase the number of FLARE blocks. To our knowledge, this is the first attention-based neural surrogate model trained on one million points on a single GPU without memory offloading or distributed computing.
4.3 Model analysis and ablations
In this section, we scale FLARE to examine how expressivity trades off against time and memory complexity. Complementary ablations in Appendix F analyze architectural choices such as the depth of residual MLPs in key/value projections, the design of the feedforward block, and the effect of head dimension on accuracy.
Time and memory complexity.
Although vanilla self-attention has the lowest memory cost thanks to the flash-attention algorithm, which eliminates the need to materialize the score matrices ($Q_{h}\cdot K_{h}^{T}$), its compute time still scales poorly with the sequence length. In contrast, the compute time for FLARE exhibits strong scaling with sequence length. Its memory requirement is marginally greater than vanilla attention due to the presence of deep residual networks for key/value projections, and due to the need to materialize $Z_{h}$, the latent sequence of $M$ tokens. As these costs are marginal compared to the SDPA operation, the curves for different $M$ values of FLARE are somewhat overlapping. Finally, the compute time for Physics Attention of Transolver (Wu et al., 2024) exhibits somewhat good scaling. However, its memory cost and compute time blow up for large slice counts due to the need for materializing the projection matrices.
Figure 5: Effect of number of blocks ($B$) and number of latent tokens ($M$) on test accuracy on the elasticity (left) and darcy (right) test cases.
Number of blocks ($B$) and latent tokens ($M$).
@@ 5 Conclusion
As transformers are the backbone of modern deep learning, we postulate that an efficient attention mechanism has several applications. We also identify potential areas for improvement: FLARE’s reliance on deep residual MLPs can introduce sequential bottlenecks and increase latency, suggesting that further speedups are possible by addressing this issue. Additional enhancements include (1) incrementally increasing the number of latent tokens during training; (2) conditioning latent tokens on time for diffusion modeling; and (3) designing decoder-only variants for auto-regressive modeling.
LLM Usage
LLMs assisted with copyediting (grammar/clarity) and with scaffolding boilerplate in our experimental scripts (e.g., debugging, refactors) inside an IDE with LLM support. LLMs did not generate ideas, architectures, analyses, results, or datasets. All LLM outputs were reviewed, edited, and tested by the authors, who take full responsibility for the content.
@@ Acknowledgments
This work was supported by the Air Force Research Laboratory under contract FA8650-21-F-5803 and PA Manufacturing Fellows Initiative. Zhang was supported in part by the National Science Foundation under grants CMMI-1953323 and CBET-2332084. The authors appreciate the support of Camfer, Inc. (Camfer, 2025) and Professor Amir Barati Farimani in providing computing resources. Part of this research was conducted on the Bridges-2 Supercomputer at the Pittsburgh Supercomputing Center, and parts were conducted using ORCHARD, a high-performance cloud computing cluster made available by Carnegie Mellon University. The authors thank Jay Pathak of Ansys, Inc. (Ansys, 2025) for insightful discussions and Andrew Porco for his assistance with data processing.
Appendix A Architecture details
A.1 Input/ output projection

### v3

@@ 3.3 Spectral analysis
The matrix $W_{h}$ in Eq. 9 represents the attention weights between $N$ tokens, where $W_{h}]_{ij}$ quantifies how much token $j$ communicates to token $i$ within head $h$. Since $W_{h}$ has rank at most $M$, the model can capture at most $M$ independent global communication patterns per head. The eigenvalues of $W$ indicate the relative importance or energy of each latent dimension in forming the attention matrix $W$. We apply an algorithm to obtain the eigendecomposition of $W$ in $\mathcal{O}(M^{3}+M^{2}N)$ time compared to $\mathcal{O}(N^{3})$ for a dense communication matrix. The algorithm is predicated on computing the eigenspectra of an $M\times M$ matrix $JJ^{T}$ where $J\in\mathbb{R}^{M\times N}$ is chosen such that $J^{T}J$ is similar to $W$. The algorithm is detailed in Section C.1, and summarized in Algorithm 1. Figure 12 presents the $M$ nonzero eigenvalues of $W_{h}$ for an FLARE model trained on the Elasticity benchmark problem with 972 points per input. The distinct spectra of the heads indicates that each head learns distinct attention patterns.
The eigenvalue analysis detailed in Section C.2 shows that while FLARE provides capacity for rank-$M$ attention, the model learns to use only a small fraction of this in early blocks indicating effective compression. In deeper blocks, more of the latent capacity is utilized, with diverse spectral profiles across heads, validating our design choice of independent head-wise projections.
4 Benchmark dataset for additive manufacturing
We introduce a benchmark dataset of laser powder bed fusion (LPBF) simulations designed to evaluate surrogate models on large, irregular 3D geometries. LPBF is a widely used metal additive manufacturing process in which a laser fuses thin layers of powder, producing complex parts but often inducing residual stresses and distortions that can lead to build failures (Zhang et al., ). To capture these effects, we simulate the thermo–mechanical LPBF process on thousands of geometries drawn from the Fusion 360 segmentation dataset (Lambourne et al., 2021).
Our benchmark task is predicting vertical ($Z$) displacement field for each geometry, a quantity directly associated with recoater-blade collisions and build-failure risk in LPBF. Each sample consists of the 3D mesh coordinates as input and the final $Z$-displacement at all nodes as output. The dataset spans a wide variety of part shapes, mesh resolutions, and deformation magnitudes, making it a challenging and practically meaningful testbed for large-scale surrogate modeling. Additional details such as statistics and visualizations are provided in Appendix H.
@@ 5.1 PDE surrogate benchmarks
We consider a diverse set of benchmark datasets (Table 3) for regressing PDE solutions on point clouds spanning structured and unstructured grids with up to 50,000 points. Note that FLARE is mesh-agnostic, and operates solely on the input point cloud. The 2D Elasticity, Darcy, Airfoil, and Pipe benchmarks (Li et al., 2020, 2023b) cover a wide range of physical phenomena, and the 3D DrivAerML benchmark (Ashton et al., 2024) provides automotive aerodynamic simulations. Visualizations of the Elasticity, Darcy, and Pipe benchmark problems are presented in Figure 4. Additional details of the dataset are presented in Section D.2. We also introduce a 3D field-prediction benchmark derived from laser powder bed fusion (LPBF) simulations, with diverse 3D-printed parts containing up to 50,000 grid points (see Appendix H).
@@
We compare FLARE with state-of-the-art PDE surrogates: generic attention models (vanilla Transformer (Vaswani et al., 2017), PerceiverIO (Jaegle et al., 2021a)); attention-based PDE surrogates (Transolver (Wu et al., 2024), LNO (Wang and Wang, 2024b)); and the neural operator GNOT (Hao et al., 2023). We exclude graph-based models because graph connectivity is unavailable for most problems. Transolver++ (Luo et al., 2025) is a concurrent work; we tested the official implementation with reported hyperparameters, but were unable to reproduce the reported numbers, so we do not include it in our comparisons (details in Section D.3). We follow the experimental setup of Transolver as it is the preeminent surrogate model and attempt to match its parameter count. Note that Transolver can be instantiated in two configurations: without convolution, where point-to-point communication relies solely on physics attention, and with convolution, where convolution layers are added to inject information from neighboring points when the input grid is structured. We evaluate these two configurations separately to isolate the impact of convolution versus physics attention. In our model, we choose not to employ any convolution layers and rely entirely on FLARE for token mixing. Because these PDE problems are relatively small (up to 50,000 points), we train all models in FP32. As shown in Figure 8, the vanilla Transformer is drastically slower than FLARE and Transolver on large point clouds; accordingly, we evaluate it only on problems up to $\sim 10{,}000$ points.
@@ Discussion.
Table 2: Accuracy (%) of different transformer models on Long Range Arena benchmark tasks (Tay et al., ). The best result (highest accuracy) is bold and the second best is underlined.
Model ListOps Text Retrieval Image Pathfinder-32 Avg
Vanilla attention 36.70 64.93 77.18 38.02 70.52 57.47
Linear attention 17.15 66.00 71.84 09.86 75.00 47.97
Linformer 36.70 53.00 64.72 41.88 70.09 53.28
Norm attention 17.10 63.08 76.07 36.94 70.15 52.67
Performer 35.90 64.21 68.42 35.36 53.83 51.54
FLARE (ours) 36.15 64.00 77.30 40.96 71.91 58.06
@@ 5.2 Field-prediction on million-point geometries
The benchmark problems in Section 5.1 contain a variety of PDE problems, but are relatively small compared to industrial use cases that demand PDE solutions on complex geometries with millions of grid points (Ashton et al., 2024). So far, attention-based surrogate models have not been able to scale to million-scale regression problems due to quadratic time and memory complexity, as illustrated in Figure 2. The FlashAttention (Dao et al., 2022) algorithm has alleviated the memory bottleneck thanks to online softmax computation; however, these methods remain impractical due to their long training times. Furthermore, SOTA models such as Transolver and LNO are not readily expressible purely in terms of SDPA calls for their dominant $\mathcal{O}(NM)$ projection/unprojection steps; in typical implementations these steps explicitly materialize $M\times N$ projection weights, which prevents using PyTorch’s fused SDPA backends (e.g., FlashAttention-style kernels) for the $\mathcal{O}(NM)$ bottleneck.
@@
We demonstrate in Figure 5 that FLARE can scale to million-scale geometries by training on the DrivAerML dataset (Ashton et al., 2024) where each mesh is subsampled to contain one million points. These calculations are performed in mixed precision on a single Nvidia H100 80GB GPU provisioned through Google Cloud Platform. We note the clear trend in Figure 5 (left) that the error consistently decreases as we increase the number of FLARE blocks. To our knowledge, this is the first attention-based neural surrogate model trained on one million points on a single GPU without memory offloading or distributed computing. Additional details are presented in Appendix E.
5.3 Long Range Arena benchmark problems
To demonstrate that FLARE is not limited to PDE surrogate modeling, we also evaluate it on the Long Range Arena (LRA) benchmark suite (Tay et al., ), which covers a diverse set of long-context tasks ranging from logical reasoning and text classification to retrieval, image classification, and visual pathfinding. While many top-performing LRA models rely on architectures specialized for 1D sequence structure (e.g., linear-attention kernels and state-space models such as MEGA (Ma et al., 2022) and Mamba (Gu and Dao, )), these methods typically assume an inherent token ordering and are therefore not directly applicable to the unordered point clouds and unstructured meshes that motivate this work. In contrast, FLARE is fully permutation-equivariant and makes no such structural assumptions. As shown in Table 2, FLARE achieves the highest average accuracy among the attention baselines we evaluate, outperforming general-purpose efficient-attention mechanisms (Linformer (Wang et al., 2020), Performer (Choromanski et al., 2020), Norm Attention (Qin et al., 2022), Linear Attention) and even vanilla self-attention on average. These results indicate that the low-rank formulation underlying FLARE provides a robust and broadly applicable inductive bias, extending beyond PDE surrogates to heterogeneous long-range tasks.
5.4 Model analysis and ablation studies
We conduct focused ablations in Appendix G that causally isolate the contribution of each component and clarify trade-offs between expressivity, time, and memory complexity. Runnable code for all experiments, benchmarks, and ablations is provided in the supplementary material.
We first examine time and memory scaling as a function of sequence length. FLARE exhibits true linear $\mathcal{O}(NM)$ scaling in both runtime and peak memory usage, even up to one million tokens. This behavior follows directly from FLARE’s SDPA-compatible formulation, which avoids materializing attention matrices or $M\times N$ projection weights and allocates sequence-length–dependent memory only for input activations.
Next, we study the trade-off between model depth and latent dimensionality. Across all PDE benchmarks, including the 1M-point DrivAerML dataset, accuracy improves consistently with more encode–decode blocks, while increasing the number of latent tokens $M$ yields diminishing returns except on inherently high-rank tasks. This demonstrates that repeated low-rank mixing through depth is a more effective driver of global communication than increasing the rank of a single attention operator.
We then evaluate head-wise parallel low-rank pathways. Varying the number of attention heads while keeping total width fixed shows that FLARE performs best with many small heads rather than a few large ones, indicating that aggregating multiple independent low-rank operators approximates richer attention patterns than concentrating capacity into fewer projections.
Next, we compare latent-space self-attention versus encode–decode blocks and find that latent self-attention consistently worsens accuracy and increases computation, whereas adding encode–decode blocks improves performance; the best results occur when latent self-attention is entirely removed. Then, we study shared versus independent latent tokens and observe that sharing latents across heads collapses the induced spectra and degrades accuracy, while independent latent slices yield diverse eigenvalue decay profiles and lower error.
Together, these ablations validate FLARE’s architectural choices: explicit low-rank operator factorization, repeated encode-decode mixing, and head-wise latent independence. Its gains, therefore, do not stem from increased parameter count or implementation-level optimizations.
@@ 6 Conclusion
As transformers are the backbone of modern deep learning, we postulate that an efficient attention mechanism has several applications. We also identify potential areas for improvement: FLARE’s reliance on deep residual MLPs can introduce sequential bottlenecks and increase latency, suggesting that further speedups are possible by addressing this issue. Additional enhancements include (1) incrementally increasing the number of latent tokens during training; (2) conditioning the latent queries/tokens on the input to relax the fixed-$Q$ constraint; (3) conditioning latent tokens on time for diffusion modeling; and (4) designing decoder-only variants for autoregressive modeling.
@@ Acknowledgments
This work was supported by the Air Force Research Laboratory under contract FA8650-21-F-5803 and PA Manufacturing Fellows Initiative. Zhang was supported in part by the National Science Foundation under grants CMMI-1953323 and CBET-2332084. The authors appreciate the support of Camfer, Inc. (Camfer, 2025) and Professor Amir Barati Farimani in providing computing resources. Part of this research was conducted on the Bridges-2 Supercomputer at the Pittsburgh Supercomputing Center, and parts were conducted using ORCHARD, a high-performance cloud computing cluster made available by Carnegie Mellon University. The authors thank Jay Pathak of Ansys, Inc. (Ansys, 2025) for insightful discussions, undergraduate researcher Andrew Porco for his assistance with data processing.
Impact Statement
Self-attention underpins many modern AI systems, but its quadratic computational and memory cost makes both training and inference expensive and difficult to scale. By making attention-based models practical on very large inputs within single-GPU workflows, FLARE lowers the barrier to deploying powerful sequence models across a wide range of applications.
Appendix A Extended related work
Criteria PerceiverIO Transolver LNO FLARE
Asymptotic evaluation cost* $2MN+BMN$ $2BMN+BM^{2}$ $M^{2} 2BMN2BMN$
Sequential enc/dec blocks
Headwise parallel enc/dec
Encoder/decoder coupling
Latent space attention
Fused attention compat.
  • *Evaluating cost of processing $N\gg M$ tokens ignoring feedforward layers (FFN are $\mathcal{O}(N)$).
  • Physics Attention in Transolver uses the same projection weights for all heads, whereas PerceiverIO and FLARE use cross-attention projection where each head learns a distinct slice of the latent tokens.
Figure 6: Comparison of (a) PerceiverIO (Jaegle et al., 2021a), (b) Transolver (Wu et al., 2024), (c) LNO (Wang and Wang, 2024a), and (d) FLARE (ours). Each model contains $B$ blocks and $M$ latents. The criteria in the table are described in detail in Appendix A.
Figure 6 provides a structural and computational comparison between PerceiverIO Jaegle et al. (2021a), Transolver Wu et al. (2024), Latent Neural Operator (LNO) Wang and Wang (2024b), and FLARE. All four architectures introduce a latent bottleneck to reduce the quadratic cost of self-attention, but they differ fundamentally in how latent tokens are used, how information is mixed across tokens, and how the resulting attention operator is realized and implemented.
Perceiver and PerceiverIO.
Perceiver and PerceiverIO (Jaegle et al., 2021b, a) introduce a general latent-attention framework in which a fixed-size latent array of length $M$ attends to an input sequence of length $N$ via cross-attention. This step maps the input into latent space, after which multiple layers of latent self-attention are applied. Optionally, PerceiverIO performs additional cross-attention from the latent array back to the input or output tokens. This design reduces the cost of attention on the input sequence and enables flexible input and output modalities.
However, Perceiver-style models use latent tokens as a computational workspace: information is repeatedly transformed within the latent space through self-attention. As a result, the effective input-input communication pattern is a deep composition of latent transformations rather than a single, explicit attention operator. While expressive, this structure makes it difficult to characterize the induced token-to-token interactions or to interpret the model as implementing a specific low-rank attention operator on the input sequence.
Transolver and Latent Neural Operator.
Transolver (Wu et al., 2024) and Latent Neural Operator (LNO) (Wang and Wang, 2024b) adapt latent-attention ideas specifically for PDE surrogate modeling on point clouds and meshes. Both architectures project variable-length inputs into a fixed-length latent representation and use latent self-attention to aggregate global information. A key difference from PerceiverIO is that Transolver performs projection and unprojection in every transformer block, allowing global context to accumulate progressively with depth. LNO, by contrast, performs a single projection-unprojection step followed by latent self-attention.
In practice, both Transolver and LNO rely on feature-projection mechanisms that explicitly construct or manipulate $M\times N$ projection weights, typically via feature expansion followed by attention or linear layers. These projection layers introduce additional memory and compute overhead that becomes significant for large $N$ and $M$. Moreover, both models rely on latent-space self-attention as the primary mechanism for global mixing, which introduces an $\mathcal{O}(M^{2})$ cost per block.
We note that Transolver can be instantiated with or without convolutional layers. Without convolution, point-to-point communication relies solely on physics attention; with convolution, local neighborhood aggregation is explicitly injected before attention. We evaluate both configurations separately to disentangle the effects of convolutional inductive bias and latent attention. Our results in Table 1 show that Transolver without convolution performs poorly, indicating that its physics-attention mechanism alone is insufficient for effective global communication. When convolutions are included, local aggregation significantly improves performance, suggesting that much of the gain arises from convolutional preprocessing rather than latent attention itself. In contrast, FLARE does not rely on any convolutional layers and achieves strong performance using attention-based token mixing alone.
FLARE: latent routing as a low-rank operator.
FLARE differs from prior latent-attention architectures in how latent tokens are used. Rather than serving as a computational workspace, latent tokens in FLARE act purely as a routing mechanism. Each block consists of exactly two cross-attention operations (encode and decode) that together induce an explicit rank-$\leq M$ linear operator on the input tokens. FLARE eliminates latent-space self-attention entirely, ensuring that all token mixing occurs in the input space through a single encode-decode factorization.
A further distinction is FLARE’s use of head-wise independent latent slices. Each attention head projects into its own latent subspace, forming multiple parallel low-rank projection-reconstruction pathways. This contrasts with Transolver, which shares projection weights across heads, and LNO, which uses a single global projection. As shown in our spectral analysis and ablations, head-wise independence enables different heads to specialize in complementary low-rank routing patterns rather than collapsing into a shared subspace.
Summary of architectural trade-offs.
As summarized in Figure 6, FLARE combines several desirable properties of prior approaches while avoiding their key limitations. Like Transolver, it performs encode-decode operations in every block, enabling progressive global mixing. Like PerceiverIO, it relies on cross-attention and is fully compatible with fused SDPA kernels. Unlike all three prior methods, FLARE eliminates latent self-attention and uses head-wise independent latent routing, yielding a simple, interpretable, and strictly linear-time attention operator. These differences explain both FLARE’s scalability to extreme sequence lengths and its strong empirical performance.
Below we briefly clarify the criteria used in Figure 6. Asymptotic evaluation cost refers to the dominant attention-related complexity for processing $N\gg M$ tokens with $B$ blocks, ignoring feed-forward layers which are linear in $N$. The remaining criteria describe architectural structure rather than cost. Sequential encode-decode blocks indicate whether global context is progressively refined across depth (as in Transolver and FLARE). Head-wise parallel encode-decode denotes whether each attention head learns an independent latent projection, which is true for PerceiverIO and FLARE but not for Transolver or LNO. Encoder/decoder coupling refers to whether encoding and decoding projections are tied, and latent space attention indicates the presence of explicit self-attention over latent tokens.
Finally, fused-attention compatibility captures whether all attention operations can be expressed using standard scaled dot-product attention (SDPA) primitives. PerceiverIO and FLARE satisfy this property, enabling efficient fused implementations, whereas Transolver and LNO rely on additional projection operations that limit hardware efficiency at large $N$. Note that Transolver can employ fused SDPA kernels for latent-space self-attention ($\mathcal{O}(M^{2})$). However, encoding and decoding ($\mathcal{O}(NM)$) operations, which form the bottleneck in these computations, are implemented with linear projections and cannot employ SDPA kernels.
Appendix B Architecture details
B.1 Input/ output projection